# Decoupling Bias, Aligning Distributions: Synergistic Fairness Optimization for Deepfake Detection

### v1

@@ 1 Introduction
Deepfake, a combination of “deep learning” and “fake”, refers to the use of deep learning techniques to replace faces in real images or videos with others, thereby generating fabricated identities. Due to their high fidelity, synthetic media can evade human perception and are maliciously exploited in financial fraud, judicial evidence manipulation, and misinformation dissemination [38, 6, 44]. The convergence of technological sophistication and social impact has led the research community to classify Deepfakes as a critical security threat demanding urgent governance in the era of artificial intelligence [7, 38, 5].
@@
To mitigate the risks posed by Deepfakes, numerous detection methods have been proposed in recent years [26, 21, 28, 43, 10]. Nevertheless, fairness has often been overlooked, resulting in inconsistent performance across demographic groups defined by gender, race, and age [20, 1]. A principal driver of these disparities is dataset imbalance: large benchmarks frequently mirror societal or cultural skews, with Caucasian faces and certain genders overrepresented (e.g., FF++ dataset [34]). Under standard empirical risk minimization, models are biased toward the majority distribution, producing unequal error rates across groups. For example, detectors commonly achieve higher accuracy on lighter than on darker skin tones [39].
Early efforts [27, 41] addressed fairness from the dataset side by balancing demographic distributions through resampling. Subsequent work [19] shifted to algorithm-level objective design and loss reweighting to reduce bias, yet generalization remained insufficient. Disentanglement was then explored to separate demographic features from domain-agnostic forgery cues, which improved fairness generalization but reduced detection accuracy [23]. Overall, existing strategies either fail to generalize or degrade detection accuracy, which motivates our approach that improves both inter- and intra-group fairness while maintaining detection performance.
@@ 2.1 Deepfake Detection
Based on prior research, deepfake detection methods can be broadly categorized into four main directions: biological feature–based detection, signal-level artifact analysis, data-driven deep neural network (DNN) classification, and forgery-trace–oriented approaches. Early studies [25] focused on extracting biological or physiological cues from facial regions to expose tampering traces. Subsequently, researchers [40] found that signal-level inconsistencies such as frequency or compression artifacts were more reliable for identifying synthetic content. Current deepfake detection research is predominantly grounded in convolutional neural network (CNN)–based classification frameworks [32, 13, 31, 8, 43], which learn discriminative forgery features through end-to-end training. As forgery techniques have evolved in complexity, recent studies have advanced toward forgery-trace–oriented approaches. For instance, Cao et al. [2] introduced a forgery detection framework that relies on reconfigurable classification learning. There are also detectors developed with large language models [18].
@@ 2.2 Stage-Wise Fairness Optimization Methods
Previous studies have documented significant demographic biases in deepfake detectors across gender, age, and race [26, 30, 42, 39, 15]. To mitigate these biases, prior work can be organized into pre-processing, in-processing, and post-processing stages [9, 23, 3].
Pre-processing methods. [27, 41] rebalance data via resampling or cross-group synthesis to alleviate distribution skew. These methods show limited adaptability to evolving generative attacks and sparsely observed intersectional combinations.
In-processing approaches. [23, 19, 24, 33, 35, 17, 9] incorporate adversarial debiasing, risk-sensitive objectives, or disentanglement modules to reduce the dependence of predictions on sensitive attributes. These mechanisms can inadvertently suppress discriminative forgery cues and degrade accuracy or robustness.
Post-processing methods. [37, 14, 29] calibrate thresholds or align output distributions conditioned on sensitive attributes. Their effectiveness is constrained by residual representation bias and instability under cross-domain shifts.
@@ 3.3 Global Distribution Alignment
DAG-FDD WACV’24 [19] Xception 1.82 4.65 94.87 96.72
DAW-FDD WACV’24 [19] Xception 0.78 9.52 95.76 97.46
PG-FDD CVPR’24 [23] Xception 0.62 4.74 96.32 97.66
Fairadapter ICASSP’25 [9] ViT-L/14 4.16 12.21 67.85 71.50
RSEF-FDD arxiv’25 [17] Xception 0.57 8.55 94.91 97.09
@@
DAG-FDD WACV’24 [19] Xception 5.48 9.20 93.43 96.72
DAW-FDD WACV’24 [19] Xception 5.43 14.69 94.15 97.46
PG-FDD CVPR’24 [23] Xception 11.13 4.78 94.52 97.66
Fairadapter ICASSP’25[9] ViT-L/14 43.22 20.39 56.03 71.50
RSEF-FDD arxiv’25 [17] Xception 8.39 5.28 93.60 97.09
@@
DAG-FDD WACV’24 [19] Xception 24.08 26.25 85.80 96.72
DAW-FDD WACV’24 [19] Xception 14.36 26.13 86.74 97.46
PG-FDD CVPR’24 [23] Xception 9.19 13.39 86.83 97.66
Fairadapter ICASSP’25 [9] ViT-L/14 86.91 42.44 45.98 71.50
RSEF-FDD arxiv’25 [17] Xception 23.64 21.74 85.77 97.09
@@ 4.1 Experimental Settings
Datasets. To evaluate the fairness generalization capability of the proposed method, we trained the model on the widely used FaceForensics++ (FF++) dataset [34] and tested it on four benchmark datasets: FF++, DeepFake Detection Challenge (DFDC) [11], DeepFakeDetection (DFD) [12], and Celeb-DF [22]. To evaluate the fairness attribute, we divided the test dataset into intra-group evaluations focusing separately on gender and race, and inter-group evaluations combining both attributes (e.g., Male-White, Female-Asian). Details of each annotated dataset are in Appendix LABEL:sec:experimental_settings.
Evaluation Metrics. For detection comparison, the Area Under Curve (AUC) is used to benchmark our approach against previous works, which aligns with the detection evaluation approach adopted in precedent works [23, 19]. Regarding fairness, we use three distinct fairness metrics to evaluate the effectiveness of our proposed method. Specifically, we report the Equal False Positive Rate ($F_{FPR}$), Demographic Parity ($F_{DP}$) and esAUCes-AUC.
@@
Baseline Methods. We consider two widely used CNN architectures to validate the effectiveness of our proposed method (i.e., Xception [4], ResNet-50 [16]). We compare our method against the latest fairness method DAG-FDD [19], DAW-FDD [19], PG-FDD [23], Fairadapter [9] and RSEF-FDD [17] in deepfake detection. The compared method also includes ‘Ori’ (a backbone with cross-entropy loss).
@@ 4.2 Experimental Results
DAG-FDD WACV’24 [19] Xception 5.29 6.68 57.41 9.26 12.51 45.91 45.30 12.92 36.42 59.13
DAW-FDD WACV’24 [19] Xception 3.60 3.67 55.88 22.36 9.72 47.77 46.80 12.69 36.63 56.90
PG-FDD CVPR’24 [23] Xception 2.35 2.57 57.71 10.10 11.49 50.53 32.16 17.71 37.42 59.64
Fairadapter ICASSP’25 [9] ViT-L/14 2.39 4.57 55.88 32.49 26.77 52.27 74.10 31.53 31.39 59.25
RSEF-FDD arxiv’25 [17] Xception 2.09 3.83 57.90 12.24 7.74 50.27 38.89 12.18 37.72 59.08
@@
DAG-FDD WACV’24 [19] Xception 8.70 14.25 66.23 19.44 13.26 62.61 23.60 13.80 63.08 72.29
DAW-FDD WACV’24 [19] Xception 8.71 12.73 68.31 14.76 7.80 66.47 24.75 13.78 59.24 74.09
PG-FDD CVPR’24 [23] Xception 7.95 16.29 69.47 23.99 12.47 67.78 23.71 13.30 62.63 74.24
Fairadapter ICASSP’25 [9] ViT-L/14 8.59 26.84 56.84 46.14 36.35 58.34 47.03 35.69 54.77 64.94
RSEF-FDD arxiv’25 [17] Xception 1.94 35.63 68.92 17.40 27.48 67.80 24.17 27.73 62.56 74.36
@@
DAG-FDD WACV’24 [19] Xception 5.41 5.60 75.87 2.56 21.34 70.14 33.00 23.02 61.09 76.29
DAW-FDD WACV’24 [19] Xception 5.30 10.63 71.68 4.32 20.16 63.24 33.81 24.72 56.01 73.71
PG-FDD CVPR’24 [23] Xception 5.05 1.86 76.37 5.79 19.31 74.81 28.00 23.93 66.32 80.70
Fairadapter ICASSP’25 [9] ViT-L/14 6.32 7.55 63.28 11.29 12.28 56.28 29.56 33.11 48.63 68.12
RSEF-FDD arxiv’25 [17] Xception 18.50 6.11 75.97 19.45 19.40 72.65 35.34 10.97 67.24 80.54
@@
DAG-FDD WACV’24 [19] ResNet-50 1.92 6.02 91.09 6.38 3.98 91.39 32.58 11.06 80.64 94.74
DAW-FDD WACV’24 [19] ResNet-50 3.25 5.29 90.13 12.46 3.29 90.92 31.07 11.03 80.98 93.96
PG-FDD CVPR’24 [23] ResNet-50 3.05 3.06 93.57 10.89 2.83 91.53 22.10 9.69 81.58 94.46
@@
DAG-FDD WACV’24 [19] ResNet-50 6.57 4.87 57.63 15.96 2.95 50.31 44.10 11.25 33.41 60.29
DAW-FDD WACV’24 [19] ResNet-50 9.12 7.14 57.22 14.60 2.81 48.94 43.31 9.55 34.71 59.78
PG-FDD CVPR’24 [23] ResNet-50 5.78 4.65 58.34 12.68 5.23 49.33 31.97 8.54 34.85 61.83
@@
DAG-FDD WACV’24 [19] ResNet-50 8.22 20.45 67.01 23.07 18.50 68.07 20.29 19.33 66.76 75.77
DAW-FDD WACV’24 [19] ResNet-50 9.03 20.85 68.33 18.6 17.65 69.14 23.84 18.55 68.74 76.51
PG-FDD CVPR’24 [23] ResNet-50 6.98 16.92 68.49 16.74 16.50 70.35 29.34 18.72 66.26 76.10
@@
DAG-FDD WACV’24 [19] ResNet-50 3.62 4.19 73.74 3.86 13.54 70.33 25.22 14.84 57.19 73.97
DAW-FDD WACV’24 [19] ResNet-50 4.81 3.90 72.12 7.36 3.21 70.57 33.45 14.30 54.52 73.58
PG-FDD CVPR’24 [23] ResNet-50 3.57 1.82 76.14 2.62 7.81 72.16 23.15 11.01 62.65 77.47
@@ 4.2.3 Performance of Different Backbone
To evaluate the fairness generalization capability of our proposed method with respect to backbone network selection, we replaced the Xception backbone with ResNet-50 [16]. As shown in Table 3, our method achieves similarly superior results across different backbone. These findings demonstrate that the proposed approach is not constrained by the choice of backbone and remains effective and applicable under various backbone settings.
@@ 4.4 Visualization Study
To better illustrate the effectiveness of our method, we visualize the Grad-CAM [36] results for Ori, DAW-FDD [19], PG-FDD [23], RSEF-FDD [17] and our method, as shown in Fig. 4. The Grad-CAM results reveal that Ori, in the absence of constraints, tends to overfit to small local regions or focus on background noise outside the facial area. Due to insufficient generalization, DAW-FDD performs similarly to Ori on cross-domain datasets. PG-FDD and RSEF-FDD boost fairness generalization and reduce local overfitting, but they are still vulnerable to non-facial background noise. In contrast, the activation maps of our method demonstrate that the model consistently attends to salient facial features across different datasets.

### v2

@@ 1 Introduction
Deepfake, a combination of “deep learning” and “fake”, refers to the use of deep learning techniques to replace faces in real images or videos with others, thereby generating fabricated identities. Due to their high fidelity, synthetic media can evade human perception and are maliciously exploited in financial fraud, judicial evidence manipulation, and misinformation dissemination [tolosana2020deepfakes, ding2021anti, zhou2025breaking]. The convergence of technological sophistication and social impact has led the research community to classify Deepfakes as a critical security threat demanding urgent governance in the era of artificial intelligence [ding2022securing, tolosana2020deepfakes, citron2019deepfakes].
@@
To mitigate the risks posed by Deepfakes, numerous detection methods have been proposed in recent years [masood2023deepfakes, leibowicz2021deepfake, passos2024review, ye2024decoupling, fan2025generating]. Nevertheless, fairness has often been overlooked, resulting in inconsistent performance across demographic groups defined by gender, race, and age [krishnan2020understanding, albiero2021gendered]. A principal driver of these disparities is dataset imbalance: large benchmarks frequently mirror societal or cultural skews, with Caucasian faces and certain genders overrepresented (e.g., FF++ dataset [rossler2019faceforensics++]). Under standard empirical risk minimization, models are biased toward the majority distribution, producing unequal error rates across groups. For example, detectors commonly achieve higher accuracy on lighter than on darker skin tones [trinh2021examination].
Early efforts [nadimpalli2022gbdf, wang2022fairness] addressed fairness from the dataset side by balancing demographic distributions through resampling. Subsequent work [ju2024improving] shifted to algorithm-level objective design and loss reweighting to reduce bias, yet generalization remained insufficient. Disentanglement was then explored to separate demographic features from domain-agnostic forgery cues, which improved fairness generalization but reduced detection accuracy [lin2024preserving]. Overall, existing strategies either fail to generalize or degrade detection accuracy, which motivates our approach that improves both inter- and intra-group fairness while maintaining detection performance.
@@ 2.1 Deepfake Detection
Based on prior research, deepfake detection methods can be broadly categorized into four main directions: biological feature–based detection, signal-level artifact analysis, data-driven deep neural network (DNN) classification, and forgery-trace–oriented approaches. Early studies [lyu2022deepfake] focused on extracting biological or physiological cues from facial regions to expose tampering traces. Subsequently, researchers [wang2022deepfake] found that signal-level inconsistencies such as frequency or compression artifacts were more reliable for identifying synthetic content. Current deepfake detection research is predominantly grounded in convolutional neural network (CNN)–based classification frameworks [ramachandran2021experimental, guo2022robust, pu2022learning, ding2024disrupting, ye2024decoupling], which learn discriminative forgery features through end-to-end training. As forgery techniques have evolved in complexity, recent studies have advanced toward forgery-trace–oriented approaches. For instance, Cao et al. [cao2022end] introduced a forgery detection framework that relies on reconfigurable classification learning. There are also detectors developed with large language models [he2025vlforgery].
@@ 2.2 Stage-Wise Fairness Optimization Methods
Previous studies have documented significant demographic biases in deepfake detectors across gender, age, and race [masood2023deepfakes, pu2022fairness, xu2022comprehensive, trinh2021examination, hazirbas2021towards]. To mitigate these biases, prior work can be organized into pre-processing, in-processing, and post-processing stages [ding2025fairadapter, lin2024preserving, cheng2025fair].
Pre-processing methods. [nadimpalli2022gbdf, wang2022fairness] rebalance data via resampling or cross-group synthesis to alleviate distribution skew. These methods show limited adaptability to evolving generative attacks and sparsely observed intersectional combinations.
In-processing approaches. [lin2024preserving, ju2024improving, little2022fairness, roh2020fr, sarhan2020fairness, he2024a3, ding2025fairadapter] incorporate adversarial debiasing, risk-sensitive objectives, or disentanglement modules to reduce the dependence of predictions on sensitive attributes. These mechanisms can inadvertently suppress discriminative forgery cues and degrade accuracy or robustness.
Post-processing methods. [tifrea2023frappe, hardt2016equality, pleiss2017fairness] calibrate thresholds or align output distributions conditioned on sensitive attributes. Their effectiveness is constrained by residual representation bias and instability under cross-domain shifts.
@@ 3.3 Global Distribution Alignment
DAG-FDD WACV’24 [ju2024improving] Xception 1.82 4.65 94.87 96.72
DAW-FDD WACV’24 [ju2024improving] Xception 0.78 9.52 95.76 97.46
PG-FDD CVPR’24 [lin2024preserving] Xception 0.62 4.74 96.32 97.66
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 4.16 12.21 67.85 71.50
RSEF-FDD arxiv’25 [he2024a3] Xception 0.57 8.55 94.91 97.09
@@
DAG-FDD WACV’24 [ju2024improving] Xception 5.48 9.20 93.43 96.72
DAW-FDD WACV’24 [ju2024improving] Xception 5.43 14.69 94.15 97.46
PG-FDD CVPR’24 [lin2024preserving] Xception 11.13 4.78 94.52 97.66
Fairadapter ICASSP’25[ding2025fairadapter] ViT-L/14 43.22 20.39 56.03 71.50
RSEF-FDD arxiv’25 [he2024a3] Xception 8.39 5.28 93.60 97.09
@@
DAG-FDD WACV’24 [ju2024improving] Xception 24.08 26.25 85.80 96.72
DAW-FDD WACV’24 [ju2024improving] Xception 14.36 26.13 86.74 97.46
PG-FDD CVPR’24 [lin2024preserving] Xception 9.19 13.39 86.83 97.66
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 86.91 42.44 45.98 71.50
RSEF-FDD arxiv’25 [he2024a3] Xception 23.64 21.74 85.77 97.09
@@ 4.1 Experimental Settings
Datasets. To evaluate the fairness generalization capability of the proposed method, we trained the model on the widely used FaceForensics++ (FF++) dataset [rossler2019faceforensics++] and tested it on four benchmark datasets: FF++, DeepFake Detection Challenge (DFDC) [r10], DeepFakeDetection (DFD) [r12], and Celeb-DF [li2020celeb]. To evaluate the fairness attribute, we divided the test dataset into intra-group evaluations focusing separately on gender and race, and inter-group evaluations combining both attributes (e.g., Male-White, Female-Asian). Details of each annotated dataset are in Appendix LABEL:sec:experimental_settings.
Evaluation Metrics. For detection comparison, the Area Under Curve (AUC) is used to benchmark our approach against previous works, which aligns with the detection evaluation approach adopted in precedent works [lin2024preserving, ju2024improving]. Regarding fairness, we use three distinct fairness metrics to evaluate the effectiveness of our proposed method. Specifically, we report the Equal False Positive Rate ($F_{FPR}$), Demographic Parity ($F_{DP}$) and esAUCes-AUC.
@@
Baseline Methods. We consider two widely used CNN architectures to validate the effectiveness of our proposed method (i.e., Xception [chollet2017xception], ResNet-50 [he2016deep]). We compare our method against the latest fairness method DAG-FDD [ju2024improving], DAW-FDD [ju2024improving], PG-FDD [lin2024preserving], Fairadapter [ding2025fairadapter] and RSEF-FDD [he2024a3] in deepfake detection. The compared method also includes ‘Ori’ (a backbone with cross-entropy loss).
@@ 4.2 Experimental Results
DAG-FDD WACV’24 [ju2024improving] Xception 5.29 6.68 57.41 9.26 12.51 45.91 45.30 12.92 36.42 59.13
DAW-FDD WACV’24 [ju2024improving] Xception 3.60 3.67 55.88 22.36 9.72 47.77 46.80 12.69 36.63 56.90
PG-FDD CVPR’24 [lin2024preserving] Xception 2.35 2.57 57.71 10.10 11.49 50.53 32.16 17.71 37.42 59.64
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 2.39 4.57 55.88 32.49 26.77 52.27 74.10 31.53 31.39 59.25
RSEF-FDD arxiv’25 [he2024a3] Xception 2.09 3.83 57.90 12.24 7.74 50.27 38.89 12.18 37.72 59.08
@@
DAG-FDD WACV’24 [ju2024improving] Xception 8.70 14.25 66.23 19.44 13.26 62.61 23.60 13.80 63.08 72.29
DAW-FDD WACV’24 [ju2024improving] Xception 8.71 12.73 68.31 14.76 7.80 66.47 24.75 13.78 59.24 74.09
PG-FDD CVPR’24 [lin2024preserving] Xception 7.95 16.29 69.47 23.99 12.47 67.78 23.71 13.30 62.63 74.24
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 8.59 26.84 56.84 46.14 36.35 58.34 47.03 35.69 54.77 64.94
RSEF-FDD arxiv’25 [he2024a3] Xception 1.94 35.63 68.92 17.40 27.48 67.80 24.17 27.73 62.56 74.36
@@
DAG-FDD WACV’24 [ju2024improving] Xception 5.41 5.60 75.87 2.56 21.34 70.14 33.00 23.02 61.09 76.29
DAW-FDD WACV’24 [ju2024improving] Xception 5.30 10.63 71.68 4.32 20.16 63.24 33.81 24.72 56.01 73.71
PG-FDD CVPR’24 [lin2024preserving] Xception 5.05 1.86 76.37 5.79 19.31 74.81 28.00 23.93 66.32 80.70
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 6.32 7.55 63.28 11.29 12.28 56.28 29.56 33.11 48.63 68.12
RSEF-FDD arxiv’25 [he2024a3] Xception 18.50 6.11 75.97 19.45 19.40 72.65 35.34 10.97 67.24 80.54
@@
DAG-FDD WACV’24 [ju2024improving] ResNet-50 1.92 6.02 91.09 6.38 3.98 91.39 32.58 11.06 80.64 94.74
DAW-FDD WACV’24 [ju2024improving] ResNet-50 3.25 5.29 90.13 12.46 3.29 90.92 31.07 11.03 80.98 93.96
PG-FDD CVPR’24 [lin2024preserving] ResNet-50 3.05 3.06 93.57 10.89 2.83 91.53 22.10 9.69 81.58 94.46
@@
DAG-FDD WACV’24 [ju2024improving] ResNet-50 6.57 4.87 57.63 15.96 2.95 50.31 44.10 11.25 33.41 60.29
DAW-FDD WACV’24 [ju2024improving] ResNet-50 9.12 7.14 57.22 14.60 2.81 48.94 43.31 9.55 34.71 59.78
PG-FDD CVPR’24 [lin2024preserving] ResNet-50 5.78 4.65 58.34 12.68 5.23 49.33 31.97 8.54 34.85 61.83
@@
DAG-FDD WACV’24 [ju2024improving] ResNet-50 8.22 20.45 67.01 23.07 18.50 68.07 20.29 19.33 66.76 75.77
DAW-FDD WACV’24 [ju2024improving] ResNet-50 9.03 20.85 68.33 18.6 17.65 69.14 23.84 18.55 68.74 76.51
PG-FDD CVPR’24 [lin2024preserving] ResNet-50 6.98 16.92 68.49 16.74 16.50 70.35 29.34 18.72 66.26 76.10
@@
DAG-FDD WACV’24 [ju2024improving] ResNet-50 3.62 4.19 73.74 3.86 13.54 70.33 25.22 14.84 57.19 73.97
DAW-FDD WACV’24 [ju2024improving] ResNet-50 4.81 3.90 72.12 7.36 3.21 70.57 33.45 14.30 54.52 73.58
PG-FDD CVPR’24 [lin2024preserving] ResNet-50 3.57 1.82 76.14 2.62 7.81 72.16 23.15 11.01 62.65 77.47
@@ 4.2.3 Performance of Different Backbone
To evaluate the fairness generalization capability of our proposed method with respect to backbone network selection, we replaced the Xception backbone with ResNet-50 [he2016deep]. As shown in Table 3, our method achieves similarly superior results across different backbone. These findings demonstrate that the proposed approach is not constrained by the choice of backbone and remains effective and applicable under various backbone settings.
@@ 4.4 Visualization Study
To better illustrate the effectiveness of our method, we visualize the Grad-CAM [selvaraju2017grad] results for Ori, DAW-FDD [ju2024improving], PG-FDD [lin2024preserving], RSEF-FDD [he2024a3] and our method, as shown in Fig. 4. The Grad-CAM results reveal that Ori, in the absence of constraints, tends to overfit to small local regions or focus on background noise outside the facial area. Due to insufficient generalization, DAW-FDD performs similarly to Ori on cross-domain datasets. PG-FDD and RSEF-FDD boost fairness generalization and reduce local overfitting, but they are still vulnerable to non-facial background noise. In contrast, the activation maps of our method demonstrate that the model consistently attends to salient facial features across different datasets.